OpenStudy (anonymous):

Solve: 4(x - 1) = 2(x + 2) + 6 Answer 28 8 7 14

OpenStudy (sasogeek):

4(x-1) = 2(x+2)+6 4x-4 = 2x+4+6 4x-2x=10+4 2x=14 x=14/2 x=7